OpenStudy (anonymous):

(8√ 6 - √ 2) + (5√ 6 - 6√ 2)

OpenStudy (anonymous):

think sqrt(6) as x, sqrt(2) as y, the () does nothing special here, so take it out direction, then do the simple addition/subtraction

OpenStudy (anonymous):

Then the first part of the answer is 13√ 6?

OpenStudy (ghazi):

\[13 \sqrt{6}-7 \sqrt{2}= 13\sqrt {3*2}-7 \sqrt{2}\]

OpenStudy (ghazi):

\[13\sqrt{3}*13 \sqrt{2}-7 \sqrt{2}\]